OpenStudy (anonymous):

An equation is shown below: 3(4x + 5) = 12 Which of the following correctly shows the beginning steps to solve this equation? Step 1: 12x + 15 = 12 Step 2: 12x = −3 Step 1: 12x + 5 = 12 Step 2: 12x = 7 Step 1: 7x + 5 = 12 Step 2: 7x = 7 Step 1: 7x + 8 = 12 Step 2: 7x = 4

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@josephinevessey

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

The first one is correct.

OpenStudy (welshfella):

@CasidyCat please explain why

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

yes or no

OpenStudy (dayakar):

example a(b+c) = 8 a*b + a*c =8 [distributive property]

OpenStudy (anonymous):

someone tell me the answer pleasee

OpenStudy (dayakar):

3(4x+5) =12 3*4x +3*5 =12 simplify it

OpenStudy (anonymous):

O_O

OpenStudy (anonymous):

alright